Determine the equation of the line that passes through point A(-1, 7) that is perpendicular to
the line 3x - y - 2 = 0 . Express your answer in standard form.

Answer by vaibhavaggarwal29(6)

Slope of the Given Line : y = mx + b ( where m = slope)
3x - y -2 = 0
y = 3x - 2
So, Slope = 3
Slope of the perpendicular line will be : -1/3
Equation of Perpendicular Line :
(y-7)=(-1/3)(x+1)
3y - 21 = -x -1
x + 3y = 20 Answer